You want to invest $14,000 and are looking for safe investment options. Your bank is offering you a certificate of deposit that pays a nominal rate of 10% that is compounded daily. What is the effective rate of return that you will earn from this investment?
Answer #1

Effective annual rate of return is calculated using the below formula:

EAR= (1+r/n)^n-1

Where r is the interest rate and n is the number of compounding periods in one year.

EAR= (1+ 0.10/365)^365 -1

        = (1 + 0.0003)^365 -1

        = 1.1052 – 1

        = 0.1052*100

        = 10.52%

Therefore, the effective rate of return is 10.52%

In case of any query, kindly comment on the solution.